# MagicDistillation: Weak-to-Strong Video Distillation for Large-Scale Portrait Few-Step Synthesis

### v1

@@ Abstract
Fine-tuning open-source large-scale VDMs for the portrait video synthesis task can result in significant improvements across multiple dimensions, such as visual quality and natural facial motion dynamics. Despite their advancements, how to achieve step distillation and reduce the substantial computational overhead of large-scale VDMs remains unexplored. To fill this gap, this paper proposes Weak-to-Strong Video Distillation (W2SVD) to mitigate both the issue of insufficient training memory and the problem of training collapse observed in vanilla DMD during the training process. Specifically, we first leverage LoRA to fine-tune the fake diffusion transformer (DiT) to address the out-of-memory issue. Then, we employ the W2S distribution matching to adjust the real DiT’s parameter, subtly shifting it toward the fake DiT’s parameter. This adjustment is achieved by utilizing the weak weight of the low-rank branch, effectively alleviate the conundrum where the video synthesized by the few-step generator deviates from the real data distribution, leading to inaccuracies in the KL divergence approximation. Additionally, we minimize the distance between the fake data distribution and the ground truth distribution to further enhance the visual quality of the synthesized videos. As experimentally demonstrated on HunyuanVideo, we obtain MagicDistill, which surpasses the standard Euler, LCM, DMD and even the 28-step standard sampling in FID/FVD and VBench in 1/4-step video synthesis. The project page is in W2SVD. Our model weight is in Magic 1-For-1.
Step Distillation, Video Diffusion Model, Large-Scale Portrait video Synthesis
@@ 1 Introduction
With the emergence of foundational models (Stability.ai, 2024; Rombach et al., 2022; Touvron et al., 2023) in the era of generative intelligence, unlocking the full potential of video diffusion models (VDMs) (Guo et al., 2023; Team, 2024) and their application to portrait video synthesis (Zakharov et al., 2020; Cui et al., 2024b) has become a prominent area of academic inquiry. The rapid progress in VDMs, combined with advancements in supporting infrastructure, has significantly elevated the capabilities of talking video synthesis, as demonstrated by a series of remarkable projects such as EMO (Sun et al., 2023) and Hallo-2 (Cui et al., 2024a). Compared to the conventional U-Net architecture (Ronneberger et al., 2015) traditionally used for video synthesis, diffusion transformers (DiTs) (Peebles & Xie, 2023) have demonstrated remarkable scalability, achieving refined temporal coherence, realistic facial motion dynamics, and improved visual quality in human-centric video synthesis tasks (Yang et al., 2024a; Jafari et al., 2024; Ki et al., 2024). Notably, further scaling of DiTs is highly likely to result in even greater performance improvements. The recent introduction of the open-source HunyuanVideo model (Kong et al., 2024), boasting over 13 billion (13b+) parameters, offers a promising avenue for synthesizing expressive talking videos through fine-tuning.
However, the primary challenges associated with the 13b+ HunyuanVideo is the significant inference overhead. For instance, generating a standard 129-frame video (i.e., $\sim\!5$s) using image-to-video (I2V) synthesis can take approximately 10 minutes on a single NVIDIA H100 GPU. Such extensive GPU latency renders the practical deployment of large-scale models virtually infeasible. A solution that has been evaluated to work to address this challenge is the use of (distillation-based) accelerated sampling (Zhai et al., 2024; Li et al., 2024a), which significantly enhances the efficiency of video synthesis by drastically reducing the number of function evaluations (NFEs). Most existing methods for step distillation in VDMs are built upon the latent consistency model (LCM) paradigm (Song et al., 2023). However, the LCM framework necessitates the incorporation of the exponential moving average (EMA) mechanism, resulting in the requirement to load three models’ parameters during training. Unfortunately, through our exploratory studies, we observed that even with the use of 16 GPUs equipped with 80GB and the zero redundancy optimizer (ZeRO) stage 3 CPU offload technique (Rajbhandari et al., 2021), the out-of-memory (OOM) issue persists and cannot be resolved.
Figure 1: For 4-step models with 900 iterations, vanilla DMD suffers from training collapse, resulting in the generation of uncontrollable video. In contrast, W2SVD demonstrates robustness and avoids such issues in portrait image animation.
Consequently, we turn our attention to an EMA-independent step distillation method known as distribution matching distillation (DMD) (Yin et al., 2024b, a). This method is expected to enable the development of the few-step HunyuanVideo model, which theoretically achieves both efficiency and effectiveness. In detail, DMD aims to estimate both the real data distribution and the fake data distribution using two distinct models with non-overlapping parameters. The generator’s parameters are then updated using the score distillation sampling (SDS) algorithm (Poole et al., 2023), leveraging information from both distributions. It is clear that the training process of vanilla DMD similarly requires three models, which can result in the same OOM issue as encountered with LCM. To address this issue without compromising the generator’s final performance, we propose adding the low-rank adaptation (LoRA) on the real model (i.e., estimate real data distribution) to implement the fake model (i.e., estimate fake data distribution), ensuring efficient training and avoiding the OOM issue.
@@
Unfortunately, we observed that employing LoRA to implement the vanilla DMD presents two major issues: 1) unstable training, which is highly prone to overfitting and leads to uncontrollable synthesized video outputs, as well as 2) insufficient visual quality, as evidenced by metrics such as VBench (Huang et al., 2024a, b), FID (Heusel et al., 2017), and FVD (Unterthiner et al., 2019) (see Sec. 4). We hypothesize that the root cause of these issues lies in the fact that, as training progresses, the real model gradually fails to adequately accommodate the inputs synthesized by the generator, leading to inaccuracies in the KL divergence approximation. To address this issue, as illustrated in Fig. 2, we propose Weak-to-Strong Video Distillation (W2SVD), a method that subtly adjusts the parameters of the real model toward those of the fake model. This adjustment enhances the real model’s ability to accommodate the few-step generator’s outputs, thereby improving training stability.
Figure 2: Overview of W2SVD. The core innovation of W2SVD lies in the integration of LoRA to facilitate weak-to-strong distribution matching, enabling more stable estimation of $\mathcal{D}_{\textrm{KL}}(p_{\textrm{fake}}\|p_{\textrm{real}})$. Furthermore, it incorporates $\mathcal{D}_{\textrm{KL}}(p_{\textrm{fake}}\|p_{\textrm{gt}})$ to further improve the visual quality of the synthesized videos.
We also introduce a KL divergence approximation between the fake data distribution and the ground truth distribution (i.e., $\mathcal{D}_{\textrm{KL}}(p_{\textrm{fake}}\|p_{\textrm{gt}})$ in Fig. 2, $\mathcal{D}_{\textrm{KL}}$ stands for the KL divergence) to further enhance both subject generalization and the visual fidelity of synthesized videos. This is an optional regularization term, which our experiments have shown to effectively achieve a balance between visual quality and motion dynamics. For the 1-step model, prior DMD frameworks (Yin et al., 2024c, a) have relied on the synthesized noise-data pairs and facilitated DMD convergence through the distillation paradigm (Liu et al., 2022). However, collecting the inputs and outputs of standard 28-step sampling on HunyuanVideo demonstrates to be an exceedingly time-intensive task. Consider this, we leverage the 4-step model produced by W2SVD to directly synthesize noise-data pairs, and seamlessly integrate this lightweight process into the training phase of W2SVD.
@@
We validate our proposed approach using benchmark datasets alongside our own collected widescreen video dataset. For the former, we conduct evaluations using both FID and FVD metrics on benchmark datasets such as VFHQ (Xie et al., 2022), HTDF (Zhang et al., 2021), and Celeb-V (Zhu et al., 2022), showcasing that W2SVD achieves a significant performance leap over vanilla DMD/DMD2 and LCM. Particularly, our 4-step model even surpasses the standard 28-step teacher in most benchmarks. For the latter, we use InternVL2-26B (Chen et al., 2024) to generate textual descriptions for reference images, thereby constructing the VBench customized for portrait video synthesis. Experimental results reveal that our 4-step model outperforms other methods in 6 out of 7 metrics, surpassing the best performance achieved by the vanilla DMD2. Moreover, it demonstrates superiority over the standard 28-step pre-trained model across the vast majority of metrics.
@@
Figure 3: Illustration of W2SVD. The terms “Reg.”, “Dis.”, and “Gen.” in this figure stand for “Regularization”, “Discriminator”, and “Generator”, respectively. W2SVD primarily leverages LoRA to facilitate the training of a large-scale VDM. The weight factors $\alpha_{\textrm{strong}}$ and $\alpha_{\textrm{weak}}$ are employed to achieve the weak-to-strong distribution matching. Furthermore, the regularization loss, which incorporates the ground truth video, helps to alleviate the overfitting problem encountered with the DMD Loss. Collectively, these strategies ensure the stability of the W2SVD training process and enhance the generative quality of the resulting few-step models.
@@ 2 Related Work
Here, we briefly review the scheduler as well as the architecture of the large-scale HunyuanVideo model, the step distillation algorithm and portrait video synthesis.
HunyuanVideo.
HunyuanVideo (Kong et al., 2024) is widely recognized as the most advanced and effective open-source VDM. Its scheduler adopts the flow-matching paradigm (Lipman et al., 2022) to model both the forward and reverse processes. To be specific, flow matching defines a mapping between (Gaussian) noises $\mathbf{x}_{1}$ from a noise distribution $p_{1}(\mathbf{x}_{1})$ to samples $\mathbf{x}_{0}$ from a real data distribution $p_{0}(\mathbf{x}_{0})$ in terms of an ordinary differential equation (ODE): $d\mathbf{x}_{t}=\mathbf{v}_{\Theta}(\mathbf{x}_{t},t)dt$, where $\mathbf{x}_{t}$ and $\mathbf{v}_{\Theta}(\cdot,\cdot)$ denote the sample at timestep $t$ on the flow and the velocity estimation model, respectively. Flow matching optimizes $\mathbf{v}_{\Theta}(\cdot,\cdot)$ by solving a simple least squares regularization problem:
@@
where $\alpha_{t}$ is the weight factor, which takes the form of logit-normal by default (Karras et al., 2022). After training, sampling is done directly using the Euler algorithm. In terms of model architecture, HunyuanVideo builds upon the DiT framework but introduces the double-stream blocks in the initial 1/3 of the model. This design allows for the integration of reference text/image with noisy video. In the latter 2/3 of the model, the single-stream blocks is employed to enable more fine-grained video modeling. Notably, HunyuanVideo’s VAE is trained from scratch, achieving a downsampling from $T\times H\times W$ to $\lfloor\frac{T}{4}\rfloor\times\lfloor\frac{H}{8}\rfloor\times\lfloor\frac{W}{8}\rfloor$, where $T$, $H$, and $W$ denote the temporal length, height, and width of the video, respectively.
Step Distillation.
Step distillation enables accelerated sampling by reducing the NFEs during inference of the diffusion model. At present, the prevailing paradigms in this domain revolve around two main approaches: “few steps aligned to many steps” and “distribution matching”. Among these, LCM stands out as the primary framework for aligning few-step models with multi-step processes. A majority of the open-source few-step VDMs are constructed using LCM training, including MCM (Zhai et al., 2024), the T2V-turbo series (Li et al., 2024b), and FastVideo (Hao-AI-Lab, 2025). The paradigm “distribution matching” has demonstrated superior performance over LCM in image synthesis, as substantiated by methods such as DMD (Yin et al., 2024b), DMD2 (Yin et al., 2024a), and SiD-LSG (Zhou et al., 2024). However, its application within video synthesis remains relatively uncommon (Yin et al., 2024c).
Portrait Video Synthesis.
The core objective of portrait video synthesis is to synthesize realistic talking-head videos by ensuring that the synthesized output aligns seamlessly with the reference image, audio, or text inputs. Previous algorithms predominantly rely on UNet as the backbone alongside diffusion models for portrait video synthesis. While these approaches achieve competitive performance, such as EMO (Sun et al., 2023), Sadtalker (Zhang et al., 2023) and Hallo (Cui et al., 2024a), they face significant challenges in scaling to longer videos and maintaining high video fidelity. This essentially stems from the lack of scalability in the UNet architecture, coupled with the insufficient length of videos used during the training process. Some recent open-source DiT-based VDMs (Yang et al., 2024b; Polyak et al., 2024) have garnered significant attention due to their ability to generate videos with high fidelity and semantic faithfulness. This motivated us to fine-tune the most advanced open-source VDM, HunyuanVideo, to establish two baseline models: one for portrait image-to-video (I2V) synthesis and another for portrait text-and-image-to-video (TI2V) synthesis. As the primary focus of this work is on accelerated sampling, we do not consider audio input at this stage.
@@ 3 Method
As illustrated in Fig. 3, we present W2SVD for accelerated sampling of large-scale portrait video synthesis. W2SVD incorporates weak-to-strong (W2S) distribution matching along with ground truth supervision. This technique aims to prevent the synthesis of uncontrollable videos and to improve visual quality in the few-step video synthesis scenario. Moreover, we develop a tailored discriminator specifically for HunyuanVideo and put forward a more efficient algorithm designed for the 1-step generator distillation.
@@ 3.1 Preliminary
$\begin{split}&\nabla_{\phi}\mathcal{L}_{\textrm{DMD}}\triangleq\mathbb{E}_{t}\nabla_{\phi}\mathcal{D}_{\textrm{KL}}\left(p_{\textrm{fake}}\|p_{\textrm{real}}\right)\approx-\mathbb{E}_{t}\int_{\epsilon}\Big{(}[\mathbf{x}_{t}\\ &-\sigma_{t}\mathbf{v}_{\Theta}^{\textrm{real}}(\mathbf{x}_{t},t)-\mathbf{x}_{t}+\sigma_{t}\mathbf{v}_{\theta}^{\textrm{fake}}(\mathbf{x}_{t},t)]\frac{\partial G_{\phi}(\epsilon)}{\partial\phi}d\epsilon\Big{)},\\ \end{split}$ (2)
where $\mathbf{x}_{t}=(1-\sigma_{t})G_{\phi}(\epsilon)+\sigma_{t}\epsilon$. Here, $\epsilon$ and $\sigma_{t}$ stand for the Gaussian noise and the noise schedule, respectively. Note that this form is the result of converting DMD from DDPM (Ho et al., 2020) to flow matching. During training, to dynamically align $p_{\textrm{fake}}$ with the output distribution of $G_{\phi}$ in real time, we minimize the following loss:
$\begin{split}\mathcal{L}_{\textrm{diffusion}}&=\mathbb{E}_{t\sim\mathcal{U}[0,1],\epsilon^{\prime}\sim\mathcal{N}(0,\mathbf{I})}\Big{[}\alpha_{t}\big{\|}\epsilon^{\prime}-G_{\phi}(\epsilon)\\ &-\mathbf{v}^{\textrm{fake}}_{\theta}(\sigma_{t}\epsilon^{\prime}+(1-\sigma_{t})G_{\phi}(\epsilon),t)\big{\|}_{2}^{2}\Big{]},\\ \end{split}$ (3)
@@
$\begin{split}&\mathcal{L}_{\textrm{dis}}=\mathbb{E}_{\epsilon,\mathbf{x}^{\textrm{gt}}}\max\left(0,1\!+\!D_{\xi}(G_{\phi}(\epsilon))\right)+\\ &\max\left(0,1\!-\!D_{\xi}(\mathbf{x}^{\textrm{gt}})\right),\quad\mathcal{L}_{\textrm{gen}}=-\mathbb{E}_{\epsilon}D_{\xi}(G_{\phi}(\epsilon)),\\ \end{split}$ (4)

### v2

# MagicDistillation: Weak-to-Strong Video Distillation for Large-Scale Few-Step Synthesis

@@ Abstract
Recently, open-source video diffusion models (VDMs), such as WanX, Magic141 and HunyuanVideo, have been scaled to over 10 billion parameters. These large-scale VDMs have demonstrated significant improvements over smaller-scale VDMs across multiple dimensions, including enhanced visual quality and more natural motion dynamics. However, these models face two major limitations: (1) High inference overhead: Large-scale VDMs require approximately 10 minutes to synthesize a 28-step video on a single H100 GPU. (2) Limited in portrait video synthesis: Models like WanX-I2V and HunyuanVideo-I2V often produce unnatural facial expressions and movements in portrait videos. To address these challenges, we propose MagicDistillation, a novel framework designed to reduce inference overhead while ensuring the generalization of VDMs for portrait video synthesis. Specifically, we primarily use sufficiently high-quality talking video to fine-tune Magic141, which is dedicated to portrait video synthesis. We then employ LoRA to effectively and efficiently fine-tune the fake DiT within the step distillation framework known as distribution matching distillation (DMD). Following this, we apply weak-to-strong (W2S) distribution matching and minimize the discrepancy between the fake data distribution and the ground truth distribution, thereby improving the visual fidelity and motion dynamics of the synthesized videos. Experimental results on portrait video synthesis demonstrate the effectiveness of MagicDistillation, as our method surpasses Euler, LCM, and DMD baselines in both FID/FVD metrics and VBench. Moreover, MagicDistillation, requiring only 4 steps, also outperforms WanX-I2V (14B) and HunyuanVideo-I2V (13B) on visualization and VBench. Our project page is https://magicdistillation.github.io/MagicDistillation/.
footnotetext: Equal contribution, Project lead, Corresponding author.
@@ 1 Introduction
As generative intelligence gradually transitions from image generation to video synthesis [27, 23, 32, 6, 31], an increasing number of high-quality large-scale video diffusion models (VDMs) have emerged. Compared to traditional VDMs built on U-Net [24] (e.g., AnimateDiff [6] and ModelScope-T2V [34]), these models, leveraging diffusion transformers (DiTs) [20], have made groundbreaking advances in the quality of synthesized videos. However, even though these VDMs adhere well to physical laws and exhibit strong generalization capabilities for video synthesis across a variety of scenarios, the primary challenges associated with those VDMs is the significant inference overhead. For instance, on an advanced GPU like H100, synthesizing a 28-step, 129-frame (i.e., $\sim\!5$s) video with flash attention3 [25] using HunyuanVideo (13B) [13], Magic141 [40] or WanX (14B) [36] takes approximately 10 minutes, rendering them completely unsuited for real-time applications.
Apart from the substantial inference GPU latency that these VDMs entail, they are still woefully insufficient in synthesizing portrait videos [45, 5], which encompass both real humans and characters, and remains notably inadequate. As illustrated in Fig. 1, WanX-I2V and HunyuanVideo-I2V demonstrate evident shortcomings in motion dynamics when it comes to portrait video synthesis. Moreover, WanX-I2V exhibits issues with inconsistency in character identity. It should be noted that, historically, there have been several algorithms specifically designed to address portrait video synthesis, such as EMO [30], Hallo-2 [4], and Hallo-3 [5]. However, these algorithms have compromised their ability to synthesize general videos.
Figure 1: Comparative visualization of synthesized videos for both real humans and characters using WanX-I2V (14B), HunyuanVideo-I2V (14B), Magic141, and MagicDistillation. Notably, WanX-I2V, HunyuanVideo-I2V, and Magic141 results are generated using 50 sampling steps, while MagicDistillation achieves comparable results with only 4 sampling steps.
Figure 2: Left: Vanilla DMD suffers from training collapse (4-step models with 900 iterations). Right: Our framework integrates LoRA for weak-to-strong distribution matching and $\mathcal{D}_{\textrm{KL}}$ constraints.
To address the significant inference delays associated with large-scale VDMs and their inability to generate high-quality videos of realistic humans, full-body figures, and stylized anime characters, this paper investigates the utilization of Magic141 [40], a model trained jointly on general video and talking video datasets, as a teacher model for step distillation [46, 14]. Step distillation significantly enhances the efficiency of video synthesis by drastically reducing the number of function evaluations (NFEs). For practical implementation considerations, particularly regarding GPU memory usage and training efficiency (details provided in Appendix D), we adopt distribution matching distillation (DMD) to carry out step distillation. To further enhance the efficiency of the distillation process, as presented in Fig. 2 (Right), we apply low-rank adaptation (LoRA) to the real model (i.e., estimating the real data distribution) in order to construct the fake model (i.e., estimating the fake data distribution).
@@
Unfortunately, we observed that employing LoRA to implement vanilla DMD presents two major issues: 1) unstable training, which is highly prone to overfitting and leads to uncontrollable synthesized video outputs (see Fig. 2 (Left)), as well as 2) insufficient visual quality, as evidenced by metrics such as VBench [10, 11], FID [8] and FVD [33] (see Sec. 4). We hypothesize that the root cause of these issues lies in the fact that, as training progresses, the real model gradually fails to adequately accommodate the inputs synthesized by the generator, leading to inaccuracies in the KL divergence approximation. To address this issue, as illustrated in Fig. 2 (Right), we propose MagicDistillation, a method that subtly adjusts the parameters of the real model toward those of the fake model. This adjustment enhances the real model’s ability to accommodate the few-step generator’s outputs, thereby improving training stability. We also introduce a KL divergence approximation between the fake data distribution and the ground truth distribution (i.e., $\mathcal{D}_{\textrm{KL}}(p_{\textrm{fake}}\|p_{\textrm{gt}})$ in Fig. 2 (Right), $\mathcal{D}_{\textrm{KL}}$ stands for the KL divergence) to further enhance both subject generalization and the visual fidelity of synthesized videos. This serves as an optional regularization term, designed to guide the data distribution of the few-step generator closer to the ground truth data distribution, which can enhance the fidelity while also improving the motion dynamics.
@@
We validate our proposed approach using benchmark datasets, our own collected widescreen video dataset and the general I2V-VBench [11]. For the first one, we conduct evaluations using both FID and FVD metrics on benchmark datasets such as VFHQ [37], HTDF [49], and Celeb-V [51], showcasing that MagicDistillation achieves a significant performance leap over vanilla DMD/DMD2 and LCM. For the second one, we construct a new VBench, a benchmark specifically designed for evaluating portrait video synthesis, using widescreen portrait images (see Appendix A). The experimental results reveal that our 4-step model outperforms other methods in 6 out of 7 metrics, surpassing the best performance achieved by vanilla DMD2. For the last one, by leveraging higher-quality talking video and general video datasets, we trained a more effective 4-step model. We compared this model against WanX-I2V (14B), HunyuanVideo-I2V (13B), and Magic141. The results of our customized VBench and general I2V-VBench [11] demonstrate that MagicDistillation outperforms these VDMs, even when they use 50 sampling steps.
@@
Figure 3: Illustration of MagicDistillation. “Reg.”, “Dis.”, and “Gen.” stand for “Regularization”, “Discriminator”, and “Generator”, respectively. MagicDistillation primarily leverages LoRA to facilitate the training of a large-scale VDM. The weight factors $\alpha_{\textrm{strong}}$ and $\alpha_{\textrm{weak}}$ are employed to achieve the W2S distribution matching. Furthermore, the regularization loss, which incorporates the ground truth video, helps to alleviate the overfitting problem encountered with the DMD Loss.
@@ 2 Related Work
Here, we briefly review the scheduler as well as the architecture of the large-scale open-source VDMs, the step distillation algorithm and portrait video synthesis.
Large-Scale Open-Source VDMs. WanX [36], Step-Video [18], HunyuanVideo [13], and Magic141 [40] are widely acknowledged as some of the most advanced and efficient open-source VDMs. Their schedulers employ the flow-matching paradigm [16] to model both the forward and reverse processes. To be specific, flow matching defines a mapping between (Gaussian) noises $\mathbf{x}_{1}$ from a noise distribution $p_{1}(\mathbf{x}_{1})$ to samples $\mathbf{x}_{0}$ from a real data distribution $p_{0}(\mathbf{x}_{0})$ in terms of an ordinary differential equation (ODE): $d\mathbf{x}_{t}=\mathbf{v}_{\Theta}(\mathbf{x}_{t},t)dt$, where $\mathbf{x}_{t}$ and $\mathbf{v}_{\Theta}(\cdot,\cdot)$ denote the sample at timestep $t$ on the flow and the velocity estimation model, respectively. Flow matching optimizes $\mathbf{v}_{\Theta}(\cdot,\cdot)$ by solving a simple least squares regularization problem:
@@
where $\alpha_{t}$ is the weight factor, which takes the form of logit-normal by default [12]. After training, sampling is done directly using the Euler algorithm. Moreover, the backbone of these VDMs are all constructed based on DiT rather than UNet, effectively increasing their scaling potential and performance ceiling. Notably, VAEs utilized by these VDMs share an identical compression ratio for video input. They downsample from $T\!\times\!H\!\times\!W$ to $\lfloor\!\frac{T}{4}\!\rfloor\!\times\!\lfloor\frac{H}{8}\!\rfloor\!\times\!\lfloor\!\frac{W}{8}\!\rfloor$, where $T$, $H$, and $W$ denote the temporal length, height, and width of the video, respectively.
Step Distillation. Step distillation enables accelerated sampling by reducing the NFEs during inference of the diffusion model. At present, the prevailing paradigms in this domain revolve around two main approaches: “few steps aligned to many steps” and “distribution matching”. Among these, LCM stands out as the primary framework for aligning few-step models with multi-step processes. A majority of the open-source few-step VDMs are constructed using LCM training, including MCM [46], the T2V-turbo series [15], and FastVideo [7]. The paradigm “distribution matching” has demonstrated superior performance over LCM in image synthesis, as substantiated by methods such as DMD [42], DMD2 [41], and SiD-LSG [50]. However, its application within video synthesis remains relatively uncommon [43].
Portrait Video Synthesis. The core objective of portrait video synthesis is to synthesize realistic talking-head videos by ensuring that the synthesized output aligns seamlessly with the reference image, audio, or text inputs. Previous algorithms predominantly rely on UNet as the backbone alongside diffusion models for portrait video synthesis. While these approaches achieve competitive performance, such as EMO [30], Sadtalker [48] and Hallo [4], they face significant challenges in scaling to longer videos and maintaining high video fidelity. This essentially stems from the lack of scalability in UNet, coupled with the insufficient length of videos used during the training process.
@@ 3 Method
As illustrated in Fig. 3, we present MagicDistillation for accelerated sampling of large-scale portrait video synthesis. MagicDistillation incorporates weak-to-strong (W2S) distribution matching along with ground truth supervision. This technique aims to prevent the synthesis of uncontrollable videos and to improve visual quality in the few-step video synthesis scenario. Moreover, we develop a tailored discriminator specifically for our pre-trained VDM Magic141 and put forward a more efficient algorithm designed for the 1-step generator distillation.
@@ 3.1 Preliminary
$\begin{split}&\nabla_{\phi}\mathcal{L}_{\textrm{DMD}}\triangleq\mathbb{E}_{t}\nabla_{\phi}\mathcal{D}_{\textrm{KL}}\left(p_{\textrm{fake}}\|p_{\textrm{real}}\right)\approx-\mathbb{E}_{t}\int_{\epsilon}\Big{(}[\mathbf{x}_{t}-\sigma_{t}\mathbf{v}_{\Theta}^{\textrm{real}}(\mathbf{x}_{t},t)-\mathbf{x}_{t}+\sigma_{t}\mathbf{v}_{\theta}^{\textrm{fake}}(\mathbf{x}_{t},t)]\frac{\partial G_{\phi}(\epsilon)}{\partial\phi}d\epsilon\Big{)},\\ \end{split}$ (2)
where $\mathbf{x}_{t}=(1-\sigma_{t})G_{\phi}(\epsilon)+\sigma_{t}\epsilon$. Here, $\epsilon$ and $\sigma_{t}$ stand for the Gaussian noise and the noise schedule, respectively. Note that this form is the result of converting DMD from DDPM [9] (w.r.t, predict noise) to flow matching (w.r.t, predict velocity). During training, to dynamically align $p_{\textrm{fake}}$ with the output distribution of $G_{\phi}$ in real time, we minimize the following loss:
$\begin{split}\mathcal{L}_{\textrm{diffusion}}&=\mathbb{E}_{t\sim\mathcal{U}[0,1],\epsilon^{\prime}\sim\mathcal{N}(0,\mathbf{I})}\Big{[}\alpha_{t}\big{\|}\epsilon^{\prime}-G_{\phi}(\epsilon)-\mathbf{v}^{\textrm{fake}}_{\theta}(\sigma_{t}\epsilon^{\prime}+(1-\sigma_{t})G_{\phi}(\epsilon),t)\big{\|}_{2}^{2}\Big{]},\\ \end{split}$ (3)
@@
$\begin{split}&\mathcal{L}_{\textrm{dis}}=\mathbb{E}_{\epsilon,\mathbf{x}^{\textrm{gt}}}\max\left(0,1\!+\!D_{\xi}(G_{\phi}(\epsilon))\right)+\max\left(0,1\!-\!D_{\xi}(\mathbf{x}^{\textrm{gt}})\right),\quad\mathcal{L}_{\textrm{gen}}=-\mathbb{E}_{\epsilon}D_{\xi}(G_{\phi}(\epsilon)),\\ \end{split}$ (4)